A 1050 kg car has a velocity of 2,65 north.The car guys the rear of a stationary truck,and the bumpers lock together.The velocity of the car truck system immediately after the collision is 0,78 north.Calculate the mass of the truck

1
Expert's answer
2022-01-18T09:36:30-0500

By conservation of momentum:


"mv=(m+M)u→M=mv/u-m=2517\text{ kg}."